The sum of the ages of 5 children is equal to the age of father. After 15 years the sum of the ages of the children will be twice the age of father. Find the age of father.
Answer by mananth(16946)

let sum of ages of 5 children = x
father's age = 5x
after 15 years
(x+15*5)=2 (5x+15)
2x+150 = 10x+30
8x=120
/8
x=15 years
father's age =5*15 = 75 years